OpenStudy (anonymous):

if (1/a)+(1/b)=10, what is the value of a+b a) 1/10 b)2/5 c) 1 d)10 e)cannot be determined

OpenStudy (anonymous):

it can't be determined:$$\frac1a+\frac1b=10\\\frac{ab}a+\frac{ab}b=10ab\\a+b=10ab$$

OpenStudy (anonymous):

@oldrin.bataku so which one? e?